# Coverage Biases in High-Resolution Satellite Imagery

### v1

@@ Abstract
Satellite imagery is increasingly used to complement traditional data collection approaches such as surveys and censuses across scientific disciplines. However, we ask: Do all places on earth benefit equally from this new wealth of information? In this study, we investigate coverage bias of major satellite constellations that provide optical satellite imagery with a ground sampling distance below 10 meters, evaluating both the future on-demand tasking opportunities as well as the availability of historic images across the globe. Specifically, forward-looking, we estimate how often different places are revisited during a window of 30 days based on the satellites’ orbital paths, thus investigating potential coverage biases caused by physical factors. We find that locations farther away from the equator are generally revisited more frequently by the constellations under study. Backward-looking, we show that historic satellite image availability – based on metadata collected from major satellite imagery providers – is influenced by socio-economic factors on the ground: less developed, less populated places have less satellite images available. Furthermore, in three small case studies on recent conflict regions in this world, namely Gaza, Sudan and Ukraine, we show that also geopolitical events play an important role in satellite image availability, hinting at underlying business model decisions. These insights lay bare that the digital dividend yielded by satellite imagery is not equally distributed across our planet.
@@ Introduction
Disparities in satellite image availability have already been explored in previous works. For example, (Lesiv et al. 2018) show that the very high resolution (VHR) satellite imagery on Google Earth and Bing Maps is unevenly global distributed throughout the world, with better coverage in the USA, Europe and India. This bias is further reflected in open-source image datasets, which, according to (Shankar et al. 2017) exhibit amerocentric and eurocentric representation, potentially impacting classifier performance in developing regions. Similarly, (Kumar and Mutanga 2018) finds that processing platforms that help end-user to large amounts of imagery such as Google Earth Engine (GEE) are mainly used by institutions in developed nations. While initiatives such as the DeepGlobe 2018 Satellite Image Understanding Challenge aim to overcome entry barriers, inequities still remain (Demir et al. 2018). Although recent advances driving down GSD with ever-higher resolutions of images, the most used satellite constellations remain Landsat and Sentinel (Kumar and Mutanga 2018). Here, the evolution of data sharing policies for environmental satellite data has seen a shift towards open access, potentially benefiting various environmental and social applications (cf. (Wulder et al. 2012; Turner et al. 2015; Olbrich 2018)). (Sudmanns et al. 2020) assess the global spatio-temporal coverage and availability of Sentinel-2 satellite imagery and find that higher revisitation rates do not necessarily lead to more cloud-free imagery available – an essential precondition for optical satellite imagery to be of further use. In contrast, (Li and Roy 2017) find that combining data from multiple satellites, such as Landsat-8 and Sentinel-2A/B, can significantly improve temporal revisit intervals and observation frequency, thus enhancing opportunities for cloud-free surface observations.
@@
In general, applications for satellite imagery are plentiful: Among others, as described by (Voigt et al. 2016) and (Rufener et al. 2024), rapid emergency mapping is extensively used to assess major disaster situations, with Asia and Europe being the most intensively monitored regions. To improve global land cover maps, initiatives like OpenEarthMap – a benchmark dataset for high-resolution land cover mapping, which consists of 2.2 million segments covering 97 regions across 6 continents – have emerged (Xia et al. 2023). Satellite imagery is also used to refine poverty mappings and demographic projections (cf. (Koebe et al. 2022; Pokhriyal and Jacques 2017)). These satellite-based earth observations have potentially significant impacts on society and policy, as discussed in (Onoda and Young 2017).
@@
This study seeks to go beyond Google Maps and Bing Maps as done by (Lesiv et al. 2018) to highlight inequitable access to information by investigating the spatial and temporal coverage biases inherent in major satellite constellations providing optical satellite imagery at resolutions with a ground sampling distance (GSD) below 10 meters, namely Planetscope, Maxar, One Atlas, Imagesat, 21at and Capella Space. While we recognize that both the Landsat as well as the Sentinel programme are essential sources of information in the remote sensing community, we do not consider them in this analysis for mainly two reasons: First, with resolutions of 10m and above, both Landsat and Sentinel support use cases that are distinct to those of high-resolution imagery providers. Second, as both Landsat and Sentinel are government-led programmes, they follow a systematic (”gota catch’em all”) capture model, irrespective of specific user demands. Thus, although we expect that physical factors such as orbital paths still influence forward-looking coverage for those two programmes, we expect that the backward-looking coverage should not be influenced by socio-economic factors.
@@
Satellites are mostly organized into constellations, which are groups of satellites working together for a set of specific purposes. The most well-known example is the GPS satellites used for positional location. Therefore, in order to estimate satellite coverage of certain locations, it is essential to examine entire constellations rather than individual satellites. We consider the following constellations in our analysis: Planet Labs operates three satellite constellations to provide imagery of Earth: Dove (GSD 3-5m), RapidEye (retired, GSD 5m) and Skysat (GSD 50cm) 111www.planet.com/our-constellations/. Maxar runs two constellations: WorldView (consisting of the satellites WorldView-1, WorldView-2, WorldView-3, GeoEye-1 (GSD 30-50cm)) and WorldView Legion (GSD 30cm)  [^2] https://www.maxar.com/maxar-intelligence/constellation. 21st Century Aerospace Technology (21AT) runs the TripleSat constellation (GSD 80cm) [^3] https://www.21at.sg/productsservices/triplesat-constellation/. The EROS constellation (GSD 70-180cm) is operated by ImageSat International [^4] https://imagesatintl.com/. Airbus groups its Pléiades, Pléiades Neo, SPOT and Vision-1 and SAR satellites into one constellation (optical satellites have a GSD 30-150cm) [^5] https://space-solutions.airbus.com/imagery/our-optical-and-radar-satellite-imagery/.
In this paper, we examine four main underlying data sources: First, Two-Line Elements (TLEs) from space-track.org to determine theoretical satellite coverage in the forward-looking part of the paper; Second, Spatiotemporal Asset Catalogs (STACs) of major satellite operators and redistributors of historic satellite imagery to assess actual image availability in the backward-looking part of the paper; Third, the global dataset of (Ton et al. 2024) on socio-economic characteristics at the subnational level to investigate the socio-economic and geophysical determinants of potential coverage biases; And fourth, the Subnational Human Development Index (SHDI) published by Global Data Lab (Smits and Permanyer 2019) to complement the global dataset by (Ton et al. 2024).
@@
STACs are an open standard for organizing and sharing geospatial data and widely used in geospatial data systems to improve accessibility and interoperability of large-scale datasets. STAC organizes geospatial data as items in form of hierarchically structured catalogs that share standardized metadata fields, which, however, can be further extended with custom fields. We retrieve metadata on available imagery from Maxar’s STAC  [^6] https://developers.maxar.com/docs/discovery/api/query-the-catalog directly and for the other satellite operators via the STAC of Up42777https://docs.up42.com/developers/api-stac as a major redistributor of satellite imagery. For Planet imagery, we retrieve only metadata statistics from their STAC API [^8] https://developers.planet.com/docs/apis/data/reference/ in order to keep the amount of data manageable.
(Ton et al. 2024) present GLOPOP-S, a synthetic global population dataset on 7.3 billion individuals and 2 billion households. This open-source dataset includes socio-economic attributes such as age, gender, income, wealth, education, and settlement type by integrating data from the Luxembourg Income Study (LIS) and the Demographic and Health Survey (DHS) program. Data and code are freely available for download [^9] https://dataverse.harvard.edu/dataset.xhtml?persistentId=doi:10.7910/DVN/KJC3RH.
The subnational Human Development Index (SHDI), curated by Global Data Lab (an independent research center at Radboud University), is a composite index that is based on the three dimensions ’education’, ’health’ and ’standard of living’. Underlying indicators include years of schooling, life expectancy at birth and gross national per capita income. For further methodological details, we refer to (Smits and Permanyer 2019) and to the website of the Global Data Lab [^10] https://globaldatalab.org/shdi/about/.
@@ Results
Dependent variable: Historic image counts
@@
From the forward-looking perspective, one would expect that the absolute latitude value of a region positively influences image availability (i.e. the closer to the equator, the less images). Our regression results confirm this hypothesis. Interestingly, there appears to be also a significant positive effect of the longitude of a region. We assume this significance is caused by some omitted variables since, for example, socio-political divides, expressed through terms such as the ’Western Hemisphere’ or the ’Western World’ may loosely be represented along the longitude, but likely correlate with important factors for image availability such as a region’s economic power. Counterintuitively, cloud coverage appears to positively and significantly affect image availability. One potential explanation is that local economic hubs are often located along coastal lines, which are more prone to cloud cover compared to inland areas.
@@
In our main regression, we exclude imagery from Planet’s dove constellation for two reasons: since Planet is also hosting by far the most historic images due to its ”gotta catch’em all” strategy (cf. Figure 4), we expect the determinants of historic image availability for Planet to be structurally different from other providers and also to dominate the reported relationships. This is supported by the results for all available historic images (incl. Planet) and for Planet-only images are provided in Table 9 and Table 8, respectively, in the Appendix. Surprisingly, we observe that by including historic Planet images, the effect of the subnational HDI and the number of households becomes insignificant and/or negative, while the effects of the remaining variables remain largely unchanged. Assuming that Planet’s dove constellation captures virtually all possible landmass images, one would expect the coefficients of the variables Cloud coverage, # of households and subnational HDI to become insignificantly different from zero, especially when looking at the Planet-only results. Possible reasons for this are multiple: First, omitted variable bias, as mentioned above, could be a source of spurious correlation. As relevant, globally available, harmonized subnational data is sparse, this phenomenon could further investigated in regional validation studies, which we put for future research. Second, the assumption of the ”gotta catch’em all” strategy of the dove constellation is overly simplistic and does not entirely reflect the actual capturing approach of Planet. However, following our analysis, Planet’s capture strategy might focus on more rural, less developed areas. While this seems to be a possible positioning in the market, a more thorough market analysis could reveal more details on the dynamics in the satellite operator sector, which is beyond the scope of this paper. Third, a linear regression model does not properly capture the complex, non-linear relationships in the data. Figure 11 in the Appendix only partially supports this argument by displaying the actual vs. predicted plot for the fully-specified linear regression. Since the scatter plot shows a slight slope and heteroscedastic behaviour, thus hinting at non-linearities in the data, we repeat the analysis using a random forest model. As Figure 12 in the Appendix shows that the random forest model is able to capture these non-linearities more accurately.
The feature importance scores from the random forest model (cf. Figure 13 in the Appendix) reflect effect sizes and significance levels from Table 3: a region’s area size is the major determinant of historic image availability and both the population count and the SHDI play a statistically significant, but minor role in determining the historic image count, thus hinting at limited structural socio-economic biases in the historic image availability: less developed, more rural places have slightly less opportunities to reap the digital dividend of remote sensing due to limited data availability.
@@ Discussion
From Table 2, we can see that available optical satellite imagery with a resolution below 10m remains behind what could potentially be captured. From a business perspective that is understandable as every capture incurs a cost (e.g. wear, downlink, online provision) and if expected profits from selling images are lower than their costs, private satellite operators may opt against offering these images in the first place. As the regression results in Table  3 show, not only weather conditions such as varying cloud coverage apparently influence this trade-off, but also regional socio-economic situations. Socio-economic status, as crudely approximated by the SHDI in our study, significantly influences historic satellite image availability and, to a limited extent, explains the differences we observe between subnational regions, holding other things equal. This implies that poorer regions generally see less demand for satellite tasking, which in turn, due to the assumed business approach of most (very) high-resolution EO providers, leads to fewer archived images. This lack of demand can have multiple reasons, e.g. poorer regions have less infrastructure that is presumably worth monitoring from space such as ports or oil storage facilities and/or they lack of regional financial resources for satellite tasking, leading to downstream unmet latent demand. While the former may not necessarily hinder the ability to reap the digital dividend from earth observation for regional development, the latter presents a significant obstacle.

### v2

@@ Abstract
 [^1] To appear in the proceedings of the ICWSM’25 Workshop on Data for the Wellbeing of Most Vulnerable (DWMV). Please cite accordingly.
Satellite imagery is increasingly used to complement traditional data collection approaches such as surveys and censuses across scientific disciplines. However, we ask: Do all places on earth benefit equally from this new wealth of information? In this study, we investigate coverage bias of major satellite constellations that provide optical satellite imagery with a ground sampling distance below 10 meters, evaluating both the future on-demand tasking opportunities as well as the availability of historic images across the globe. Specifically, forward-looking, we estimate how often different places are revisited during a window of 30 days based on the satellites’ orbital paths, thus investigating potential coverage biases caused by physical factors. We find that locations farther away from the equator are generally revisited more frequently by the constellations under study. Backward-looking, we show that historic satellite image availability – based on metadata collected from major satellite imagery providers – is influenced by socio-economic factors on the ground: less developed, less populated places have less satellite images available, as demonstrated by a Gini coefficient of $0.64$. Furthermore, in three small case studies on recent conflict regions in this world, namely Gaza, Sudan and Ukraine, we show that also geopolitical events play an important role in satellite image availability, hinting at underlying business model decisions. These insights lay bare that the digital dividend yielded by satellite imagery is not equally distributed across our planet.
@@ Introduction
Disparities in satellite image availability have already been explored in previous works. For example, Lesiv et al. (2018) show that the very high resolution (VHR) satellite imagery on Google Earth and Bing Maps is unevenly global distributed throughout the world, with better coverage in the USA, Europe and India. This bias is further reflected in open-source image datasets, which, according to Shankar et al. (2017) exhibit amerocentric and eurocentric representation, potentially impacting classifier performance in developing regions. Similarly, Kumar and Mutanga (2018) finds that processing platforms that help end-user to large amounts of imagery such as Google Earth Engine (GEE) are mainly used by institutions in developed nations. While initiatives such as the DeepGlobe 2018 Satellite Image Understanding Challenge aim to overcome entry barriers, inequities still remain (Demir et al. 2018). Although recent advances driving down GSD with ever-higher resolutions of images, the most used satellite constellations remain Landsat and Sentinel (Kumar and Mutanga 2018). Here, the evolution of data sharing policies for environmental satellite data has seen a shift towards open access, potentially benefiting various environmental and social applications (cf. Wulder et al. (2012); Turner et al. (2015); Olbrich (2018)). Sudmanns et al. (2020) assess the global spatio-temporal coverage and availability of Sentinel-2 satellite imagery and find that higher revisitation rates do not necessarily lead to more cloud-free imagery available – an essential precondition for optical satellite imagery to be of further use. In contrast, Li and Roy (2017) find that combining data from multiple satellites, such as Landsat-8 and Sentinel-2A/B, can significantly improve temporal revisit intervals and observation frequency, thus enhancing opportunities for cloud-free surface observations.
@@
In general, applications for satellite imagery are plentiful: Among others, as described by Voigt et al. (2016) and Rufener et al. (2024), rapid emergency mapping is extensively used to assess major disaster situations, with Asia and Europe being the most intensively monitored regions. To improve global land cover maps, initiatives like OpenEarthMap – a benchmark dataset for high-resolution land cover mapping, which consists of 2.2 million segments covering 97 regions across 6 continents – have emerged (Xia et al. 2023). Satellite imagery is also used to refine poverty mappings and demographic projections (cf. Koebe et al. (2022); Pokhriyal and Jacques (2017)). These satellite-based earth observations have potentially significant impacts on society and policy, as discussed in Onoda and Young (2017).
@@
This study seeks to go beyond Google Maps and Bing Maps as done by Lesiv et al. (2018) to highlight inequitable access to information by investigating the spatial and temporal coverage biases inherent in major satellite constellations providing optical satellite imagery at resolutions with a ground sampling distance (GSD) below 10 meters, namely Planetscope, Maxar, One Atlas, Imagesat, 21at and Capella Space. While we recognize that both the Landsat as well as the Sentinel programme are essential sources of information in the remote sensing community, we do not consider them in this analysis for mainly two reasons: First, with resolutions of 10m and above, both Landsat and Sentinel support use cases that are distinct to those of high-resolution imagery providers. Second, as both Landsat and Sentinel are government-led programmes, they follow a systematic (”gotta catch’em all”) capture model, irrespective of specific user demands. Thus, although we expect that physical factors such as orbital paths still influence forward-looking coverage for those two programmes, we expect that the backward-looking coverage should not be influenced by socio-economic factors.
@@
Satellites are mostly organized into constellations, which are groups of satellites working together for a set of specific purposes. The most well-known example is the GPS satellites used for positional location. Therefore, in order to estimate satellite coverage of certain locations, it is essential to examine entire constellations rather than individual satellites. We consider the following constellations in our analysis: Planet Labs operates three satellite constellations to provide imagery of Earth: Dove (GSD 3-5m), RapidEye (retired, GSD 5m) and Skysat (GSD 50cm)222www.planet.com/our-constellations/. Maxar runs two constellations: WorldView (consisting of the satellites WorldView-1, WorldView-2, WorldView-3, GeoEye-1 with GSDs between 30-50cm) and WorldView Legion (GSD 30cm) [^3] https://www.maxar.com/maxar-intelligence/constellation. 21st Century Aerospace Technology (21AT) runs the TripleSat constellation (GSD 80cm) [^4] https://www.21at.sg/productsservices/triplesat-constellation/. The EROS constellation (GSD 70-180cm) is operated by ImageSat International [^5] https://imagesatintl.com/. Airbus groups its Pléiades, Pléiades Neo, SPOT and Vision-1 and SAR satellites into one constellation (optical satellites have a GSD 30-150cm) [^6] https://space-solutions.airbus.com/imagery/our-optical-and-radar-satellite-imagery/.
In this paper, we examine four main underlying data sources: First, Two-Line Elements (TLEs) from space-track.org to determine theoretical satellite coverage in the forward-looking part of the paper; Second, Spatiotemporal Asset Catalogs (STACs) of major satellite operators and redistributors of historic satellite imagery to assess actual image availability in the backward-looking part of the paper; Third, the global dataset of Ton et al. (2024) on socio-economic characteristics at the subnational level to investigate the socio-economic and geophysical determinants of potential coverage biases; And fourth, the Subnational Human Development Index (SHDI) published by Global Data Lab (Smits and Permanyer 2019) to complement the global dataset by Ton et al. (2024).
@@
STACs are an open standard for organizing and sharing geospatial data and widely used in geospatial data systems to improve accessibility and interoperability of large-scale datasets. STAC organizes geospatial data as items in form of hierarchically structured catalogs that share standardized metadata fields, which, however, can be further extended with custom fields. We retrieve metadata on available imagery from Maxar’s STAC  [^7] https://developers.maxar.com/docs/discovery/api/query-the-catalog directly and for the other satellite operators via the STAC of Up42888https://docs.up42.com/developers/api-stac as a major redistributor of satellite imagery. For Planet imagery, we retrieve only metadata statistics from their STAC API [^9] https://developers.planet.com/docs/apis/data/reference/ in order to keep the amount of data manageable.
Ton et al. (2024) present GLOPOP-S, a synthetic global population dataset on 7.3 billion individuals and 2 billion households. This open-source dataset includes socio-economic attributes such as age, gender, income, wealth, education, and settlement type by integrating data from the Luxembourg Income Study (LIS) and the Demographic and Health Survey (DHS) program. Data and code are freely available for download [^10] https://dataverse.harvard.edu/dataset.xhtml?persistentId=doi:10.7910/DVN/KJC3RH.
The subnational Human Development Index (SHDI), curated by Global Data Lab (an independent research center at Radboud University), is a composite index that is based on the three dimensions ’education’, ’health’ and ’standard of living’. Underlying indicators include years of schooling, life expectancy at birth and gross national per capita income. For further methodological details, we refer to Smits and Permanyer (2019) and to the website of the Global Data Lab [^11] https://globaldatalab.org/shdi/about/.
@@
In the last step of the backward-looking analysis, we compare the results of the regression with the Gini coefficient – a measure of inequality that, in our setup, indicates how equally distributed historic satellite images are across the different regions in the world, sorted by their respective Subnational HDI.
@@ Results
Dependent variable: Historic image count
@@
From the forward-looking perspective, one would expect that the absolute latitude value of a region positively influences image availability (i.e. the closer to the equator, the less images). Our regression results confirm this hypothesis. Interestingly, there appears to be also a significant positive effect of the longitude of a region. We assume this significance is caused by some omitted variables since, for example, socio-political divides, expressed through terms such as the ’Western Hemisphere’ or the ’Western World’ may loosely be represented along the longitude, but likely correlate with important factors for image availability such as a region’s economic power. Counterintuitively, cloud coverage appears to positively and significantly affect image availability. One potential explanation is that local economic hubs are often located along coastal lines, which are more prone to cloud cover compared to inland areas, however, further investigations into the potential underlying causal mechanisms are needed.
@@
To explore the robustness of the observed relationships, we conducted multiple ablation studies and present the results in the Appendix. In our main regression, we exclude imagery from Planet’s dove constellation for two reasons: since Planet is also hosting by far the most historic images due to its ”gotta catch’em all” strategy (cf. Figure 4), we expect the determinants of historic image availability for Planet to be structurally different from other providers and also to dominate the reported relationships. This is supported by the results for all available historic images (incl. Planet) and for Planet-only images are provided in Table 9 and Table 8, respectively, in the Appendix. Surprisingly, we observe that by including historic Planet images, the effect of the subnational HDI and the number of households becomes insignificant and/or negative, while the effects of the remaining variables remain largely unchanged. Assuming that Planet’s dove constellation captures virtually all possible landmass images, one would expect the coefficients of the variables Cloud coverage, # of households and subnational HDI to become insignificantly different from zero, especially when looking at the Planet-only results. Possible reasons for this are multiple: First, omitted variable bias, as mentioned above, could be a source of spurious correlation. As relevant, globally available, harmonized subnational data is sparse, this phenomenon could further investigated in regional validation studies, which we put for future research. Second, the assumption of the ”gotta catch’em all” strategy of the dove constellation is overly simplistic and does not entirely reflect the actual capturing approach of Planet. However, following our analysis, Planet’s capture strategy might focus on more rural, less developed areas. While this seems to be a possible positioning in the market, a more thorough market analysis could reveal more details on the dynamics in the satellite operator sector, which is beyond the scope of this paper. Third, a linear regression model does not properly capture the complex, non-linear relationships in the data. Figure 11 in the Appendix only partially supports this argument by displaying the actual vs. predicted plot for the fully-specified linear regression. Since the scatter plot shows a slight slope and heteroscedastic behaviour, thus hinting at non-linearities in the data, we repeat the analysis using a random forest model. As Figure 12 in the Appendix shows that the random forest model is able to capture these non-linearities more accurately. The feature importance scores from the random forest model (cf. Figure 13 in the Appendix) reflect effect sizes and significance levels from Table 3: a region’s area size is the major determinant of historic image availability and both the population count and the SHDI play a statistically significant, but minor role in determining the historic image count, thus hinting at limited structural socio-economic biases in the historic image availability: less developed, more rural places have slightly less opportunities to reap the digital dividend of remote sensing due to limited data availability.
In addition, as unmodelled national-level effects such as policies on restricting the capture of high-resolution imagery, the presence of domestic satellite programs or varying success of sales efforts also potentially play a role in historic satellite image availability, we account for that by re-running our main regression with country-level fixed effects (see Table 10 in the Appendix). By looking at the $R^{2}$ of the first regression, we observe that country-level fixed effects explain about 50% of the sub-national variation in historic image availability in regression 1. However, when adding our main control variable – Area Size – the additional explanatory power of the country-level fixed effects reduces substantially, implying that much of the fixed effects are related to the size and thus the size of the subnational areas of a country. Furthermore, replacing the composite index SHDI with an indicator more related to economic power of a country – the Income Index published by GDL – or running the main regression on very-high resolution imagery (i.e. $<=0.5m$) only does not change the direction and significance of the effect (cf. with Tables 11 and 12 in the Appendix). It can be noted though that the effect size more than doubles for the latter case, hinting at a stronger economic bias for very-high resolution data.
The Gini coefficient for the number of historic satellite images per $\text{km}^{2}$ across the 1726 subnational regions under study sorted by their respective SHDI is $0.64$ and therefore hints at large-scale inequalities in the distribution of available high-resolution satellite images: approx. 50% of these images are captured in the 10% of regions with the highest SHDI (cf. the Lorenz curve in Figure 14 in the Appendix). Conversely, the 40% of regions with the lowest SHDI are represented on just 10% of all available images. Interestingly, the Gini coefficient implies a stronger economic bias than the main regression shown in Table 3 above. Reasons for this can be that the Gini coefficient reflects overall distributional inequality, not how well one variable linearly explains another. In other words, if the SHDI explains which regions tend to have more images but not how many exactly, the $\text{R}^{2}$ can still be comparatively low.
@@ Discussion
From Table 2, we can see that available optical satellite imagery with a resolution below 10m remains behind what could potentially be captured. From a business perspective that is understandable as every capture incurs a cost (e.g. wear, downlink, online provision) and if expected profits from selling images are lower than their costs, private satellite operators may opt against offering these images in the first place. As the Gini coefficient and regression results in Table 3 show, not only weather conditions such as varying cloud coverage apparently influence this trade-off, but also regional socio-economic situations. Socio-economic status, as crudely approximated by the SHDI in our study, significantly influences historic satellite image availability and, to a limited extent, explains the differences we observe between subnational regions, holding other things equal. This implies that poorer regions generally see less demand for satellite tasking, which in turn, due to the assumed business approach of most (very) high-resolution EO providers, leads to fewer archived images. This lack of demand can have multiple reasons, e.g. poorer regions have less infrastructure that is presumably worth monitoring from space such as ports or oil storage facilities and/or they lack of regional financial resources for satellite tasking, leading to downstream unmet latent demand. While the former may not necessarily hinder the ability to reap the digital dividend from earth observation for regional development, the latter presents a significant obstacle.